From GlueXWiki
Jump to: navigation, search

July 21, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
    • Status
    • HV tests: results, testing procedures (Lubomir)
    • Other: wire deadening, iron source
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126
    • Tests with external tracking (Beni)
  5. Other


Participants: Eugene, Bill, Dave, Chris, Beni, Simon, and Lubomir.


- Dave (see the construction tracking): Continued with stringing wire frame#4. Wire frame#3 is waiting for soldering second phase component (Chris will do it together with Tina). Cell#5: one cathode ready, working on the second (type 2) cathode. Cell#6: working on the first cathode.

- Waiting for the chemicals for the wire deadening; probably will start testing the procedure beginning of next week. Bill: what is the wire diameter needed to deaden the wire? Lubomir: according to Garfield if we make sense wires 40um signal drops by factor of 40, for 60um - factor of 400, and for 80um -factor of 2,700. After some discussions, Bill will look what is the sagging effect of thickening the wires and Lubomir will estimate the suppression factor that we need.

- The PR for the iron source was submitted waiting for the RadCon signature, probably at the beginning of the next week.

- Investigating the sources of the high leakage currents; applied HV on the signal side of a wire frame without wires. In some of the measurements we found the current is too high because of the bad connection between the Cu tape (on which HV is applied) and the wire pads. The lowest current we got is ~10nA for one board when the capacitors are covered with Humiseal, and close to 0nA when capacitors were removed cleaned and then soldered without flux, latter done by Anatoly who put solder on the capacitors and pads, cleaned the flux and then solder the capacitors just by heating. According to Chris this is not a certified way for soldering because of possible oxyde layer.

- Conclusions for the HV problem: the capacitors itself are not the source of the leakage. Some small amount of flux found below some of the capacitors indicate that this might be a problem but this effect is reduced by covering it with humiseal that actually penetrates also below the capacitors. The humiseal helps also to cover possible sharp places on the capacitor plates that otherwise in gas would produce sparks. Lubomir proposed a procedure for testing the wire frames just before stringing. If all the leakage current of the caps is less than few tens nA the caps will be covered with humiseal, otherwise the caps showing big currents will be removed, the place cleaned and then the caps will be soldered back.

- Eugene: we can't continue with the wire frame production until it is proven the chamber can hold HV for several weeks. Also, using humiseal is allowed only if it has been used already in drift chambers. If not, we have to do our own tests including irradiation. According to Fernando humiseal has been used in Hall B chambers but outside of the gas volume. Our prototype has a lot of humiseal inside but it has not been tested with radiation. Humiseal is also rated by NASA as low outgassing.

- We decided first to investigate if there are examples of using humiseal in drift chambers. If not, we will have to do our own tests irradiating the chamber but unfortunately the accelerator is now down. We will look also for other ways to solve the HV problem. For that we will use different wire frames that we have now: some are covered with humiseal, some were cleaned and then soldered without flux.


- Bill will be working on the spacer between packages, but nos is busy with the tagger. Bill made the wire frame strong-back flatter: from +/-35/1000" now it's down to +/-11/1000".


- Chris: within two weeks we will have all the signal cables. Chris is constantly testing them: so far only one failure.

Chamber testing

- Beni showed some preliminary results (pgae 570 in the FDC logbook) from testing the first cell using external tracking. The drift time distribution doesn't look like the one for 90/10 Ar/CO2 gas mxiture but most like 80/20 and this is due to the very low CO2 rate that can't be controlled accurately. Therefore, Beni will do tests now with 40/60.